Solve the equation using the substitution method
6x+2y=16
-2x+y=-2

Answer by yhtong(11)

your questions are:
6x+2y=16 ------- 1
-2x+y=-2 ------- 2
pick either equation 1 or 2, i would choose equation 2 because its easier.
in equation 2 i would like to solve y. In this case, bring all numbers and unknown except y to the right hand side (RHS)
y=-2+2x ------ 3(i will name this equation 3)
next, substitute y in equation 3 to equation 1. That makes
6x+2(-2+2x)=16. The equation in the brackets is the value of y in equation 3 that i chose earlier
That makes
6x-4+4x=16, solve this by using linear equation
=6x+4x-4=16
=10x-4=16
=10x =16+4
=10x =20
=x =20/10
=x =2
now that you know the value of x is 2, substitute the x into equation 3 to know what value is y
y=-2+2x
y=-2+2(2)
y=-2+4
y=2
now we have both x and y values and the values are
x=2,y=2